Last year Dick was twice as old as Jane. Four years ago the sum of Dick's age and Jane's age was twice Jane's age now. How old are Dick and Jane?
Answer by stanbon(75887)

Last year Dick was twice as old as Jane. Four years ago the sum of Dick's age and Jane's age was twice Jane's age now. How old are Dick and Jane?
---------------------
Equations:
d-1 = 2(j-1)
(d-4)+(j-4) = 2j
---------------------
Rearrange::
d - 2j = -1
d - j = 8
-------
Subtract and solve for "j":
j = 9 yrs (Jane's age now)
-----------------
Solve for "d"::
d - 9 = 8
d = 17 yrs (Dick's age now)
---------------------------------
Cheers,
Stan H.